OpenStudy (anonymous):

Is there a way that you can underline words on openstudy when you ask a question or reply??

OpenStudy (anonymous):

@SolomonZelman can u pleas help me??

OpenStudy (solomonzelman):

`\underline{ text } `

OpenStudy (anonymous):

is that the only way??

OpenStudy (solomonzelman):

For example `\(\normalsize\color{blue}{ \underline{I~~like~~underlining!} }\)` would show \(\normalsize\color{blue}{ \underline{I~~like~~underlining!} }\)

OpenStudy (anonymous):

\underline{ word }

OpenStudy (anonymous):

it is not working

OpenStudy (solomonzelman):

Click equation editor, and type it there.

OpenStudy (anonymous):

\[\underline{ \text } \]

OpenStudy (anonymous):

am i doing something wrong??

OpenStudy (solomonzelman):

\[\underline{see}\] `\underline{see}`

OpenStudy (solomonzelman):

don't do `\text` ina next cell, it won't work, but type any other word, like ice-cream

OpenStudy (solomonzelman):

`\underline{see}` gives \[\underline{see}\]

OpenStudy (solomonzelman):

Repeat after me, but with a different word

OpenStudy (anonymous):

\[\underline{ ice-cream } \]

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

YAAAA it worked!!

OpenStudy (anonymous):

Thank YOu sooo much I needed that!!

OpenStudy (solomonzelman):

\(\large\color{ white }{\large {\bbox[5pt, blue ,border:2px solid white ]{ \large\text { }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\rm{c~~~units~~~~} \\ \hline \\f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red}{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

You will get to this, \(\normalsize\color{ blue }{\normalsize {\bbox[5pt, cyan ,border:2px solid purple ]{ \rm LATELY }}}\)

OpenStudy (anonymous):

wow that is soo cool

OpenStudy (solomonzelman):

I am just combining some different combinations of chart latex box, colors and fonts, and other mathematical functions.

OpenStudy (solomonzelman):

\(\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid purple ]{ ✍ }}}\)

OpenStudy (anonymous):

I dont think I will ever get that good I will just have to stick to underlining words...